# Simulating Three-dimensional Turbulence with Physics-informed Neural Networks

### v1

@@ Abstract
Turbulent fluid flows are among the most computationally demanding problems in science, requiring enormous computational resources that become prohibitive at high flow speeds. Physics-informed neural networks (PINNs) represent a radically different approach that trains neural networks directly from physical equations rather than data, offering the potential for continuous, mesh-free solutions. Here we show that appropriately designed PINNs can successfully simulate fully turbulent flows in both two and three dimensions, directly learning solutions to the fundamental fluid equations without traditional computational grids or training data. Our approach combines several algorithmic innovations including adaptive network architectures, causal training, and advanced optimization methods to overcome the inherent challenges of learning chaotic dynamics. Through rigorous validation on challenging turbulence problems, we demonstrate that PINNs accurately reproduce key flow statistics including energy spectra, kinetic energy, enstrophy, and Reynolds stresses. Our results demonstrate that neural equation solvers can handle complex chaotic systems, opening new possibilities for continuous turbulence modeling that transcends traditional computational limitations.
@@ Results
on a triply periodic domain $[0,2\pi]^{3}$. The flow evolves through vortex stretching, roll-up, and eventual breakdown, leading to a fully developed turbulent state. Our objective is to simulate the flow evolution up to final time $T=10$.
@@
Building on our successful simulation of the TGV problem, we proceed to assess the framework’s performance in three-dimensional turbulent channel flow. This prototypical wall-bounded system represents a significantly more stringent test for numerical methods. Unlike the periodic or isotropic flows previously considered, channel flow is characterized by sharp gradients near solid boundaries, inherent anisotropy, and nontrivial momentum transport. For a rigorous evaluation of predicted turbulence statistics, we specifically focus on the widely adopted benchmark case at a friction Reynolds number of $\mathrm{Re}_{\tau}=550$, for which extensive DNS statistics data are available [42, 43].
@@
We simulate incompressible channel flow within a computational domain adopting the standard configuration: $L_{x}=2\pi$ in the streamwise direction, $L_{z}=\pi$ in the spanwise direction, and a wall-normal height of $L_{y}=2$. No-slip boundary conditions are applied at the channel walls ($y=\pm 1$), with periodicity enforced in the streamwise ($x$) and spanwise ($z$) directions. The flow is driven by a constant pressure gradient of $dp/dx=0.0025$, with the kinematic viscosity set to $\nu=9.09\times 10^{-5}$. Importantly, our PINN simulation uses a snapshot of a turbulent DNS field as the initial condition. Our goal is to examine if PINNs can maintain and accurately reproduce the established turbulence statistics for this system over the time interval $t\in[0,10]$.
@@
Figure 4 shows the framework’s performance on the turbulent channel flow benchmark, compared against the reference DNS data[42, 43]. Figure 4a displays a representative instantaneous snapshot of the predicted velocity norm, which clearly captures the characteristic near-wall streaks and large-scale turbulent structures inherent to wall-bounded turbulence. Beyond this qualitative assessment, we quantify the learned statistics against reference DNS data from [42, 43].
Figure 4b shows the mean streamwise velocity profile $U^{+}$, normalized in wall units. The PINN solution closely follows the DNS data across the viscous sublayer, buffer layer, and logarithmic region, accurately reproducing the canonical mean flow. Figures 4c-4e report the root-mean-square (RMS) profiles of the velocity fluctuations $\sqrt{\left\langle u^{\prime 2}\right\rangle},\sqrt{\left\langle v^{\prime 2}\right\rangle}$, and $\sqrt{\left\langle w^{\prime 2}\right\rangle}$, all normalized by the friction velocity $u_{\tau}$. These profiles match with DNS in both peak locations and magnitudes, indicating the PINN’s ability to capture the energy distribution across different components of the turbulent fluctuations. Finally, Figures 4f-4i present the Reynolds stress components ( ${\overline{u^{\prime}u^{\prime}}}^{+},{\overline{v^{\prime}v^{\prime}}}^{+},{\overline{w^{\prime}w^{\prime}}}^{+}$, and ${\overline{u^{\prime}v^{\prime}}}^{+}$). While some minor discrepancy is observed, the predicted profiles achieve good agreement with the DNS data. We attribute this discrepancy mostly to a slight difference in the computed friction Reynolds number of our simulation ($\operatorname{Re}_{\tau}=540.3$ ) compared to the reference DNS data ($\operatorname{Re}_{\tau}=$ 546.8).
@@
Figure 4: Turbulent channel flow ($\text{Re}_{\tau}\sim 550$). (a) Instantaneous snapshot of the predicted velocity norm, illustrating characteristic near-wall streaks and large-scale turbulent structures. (b) Comparison of the mean streamwise velocity profile ($U^{+}$) between PINN predictions and DNS data, normalized in wall units. (c-e) Comparison of RMS profiles of velocity fluctuations ($u^{+}_{rms},v^{+}_{rms},w^{+}_{rms}$) between PINN predictions and DNS data, normalized by the respective friction velocity $u_{\tau}$. (f-i) Comparison of Reynolds stress components ($\overline{u^{\prime}u^{\prime}}^{+}$, $\overline{v^{\prime}v^{\prime}}^{+}$, $\overline{w^{\prime}w^{\prime}}^{+}$, and $\overline{u^{\prime}v^{\prime}}^{+}$) between PINN predictions and DNS data, normalized by the repsetive friction velocity $u_{\tau}$.
@@ Methods
$\mathcal{L}(\theta)=\underbrace{\int_{\Omega}\left|\mathcal{I}\left[\mathbf{u}_{\theta}\right](\mathbf{x})\right|^{2}\mathrm{\leavevmode\nobreak\ d}\mathbf{x}}_{\mathcal{L}_{\mathrm{ic}}(\theta)}+\underbrace{\int_{0}^{T}\int_{\partial\Omega}\left|\mathcal{B}\left[\mathbf{u}_{\theta}\right](t,\mathbf{x})\right|^{2}\mathrm{\leavevmode\nobreak\ d}\mathbf{x}\mathrm{\leavevmode\nobreak\ d}t}_{\mathcal{L}_{\mathrm{bc}}(\theta)}+\underbrace{\int_{0}^{T}\int_{\Omega}\left|\mathcal{R}\left[\mathbf{u}_{\theta}\right](t,\mathbf{x})\right|^{2}\mathrm{\leavevmode\nobreak\ d}\mathbf{x}\mathrm{\leavevmode\nobreak\ d}t}_{\mathcal{L}_{\mathrm{pde}}(\theta)},$ (6)
where $\mathcal{I}[\cdot],\mathcal{B}[\cdot]$, and $\mathcal{R}[\cdot]$ denote the initial condition, boundary condition, and Navier-Stokes residual operators, respectively. Correspondingly, $\mathcal{L}_{\text{ic }}$ penalizes violations of the initial condition, $\mathcal{L}_{\text{bc }}$ enforces the boundary conditions, and $\mathcal{L}_{\text{pde }}$ enforces the Navier-Stokes equations.
In practice, the network parameters are optimized using stochastic gradient descent and these continuous integrals are approximated using Monte Carlo sampling. Specifically, at each training iteration, collocation points are randomly sampled from the respective domains: the initial condition loss $\mathcal{L}_{\text{ic }}$ is computed from samples $\left\{\mathbf{x}_{i}^{\text{ic }}\right\}_{i=1}^{N_{\text{ic }}}\subset\Omega$, the boundary loss $\mathcal{L}_{\text{bc }}$ from points $\left\{\left(t_{i}^{\text{bc }},\mathbf{x}_{i}^{\text{bc }}\right)\right\}_{i=1}^{N_{\text{bc }}}\subset[0,T]\times\partial\Omega$, and the PDE residual loss $\mathcal{L}_{\text{pde }}$ from interior collocation points $\left\{\left(t_{i}^{\text{pde }},\mathbf{x}_{i}^{\text{pde }}\right)\right\}_{i=1}^{N_{\text{pde }}}\subset[0,T]\times\Omega$. The total loss is computed as the empirical average over these sample points,
@@
$\mathcal{L}_{\text{pde }}(\theta)=\int_{0}^{T}w(t)\int_{\Omega}\left|\mathcal{R}\left[\mathbf{u}_{\theta}\right](t,\mathbf{x})\right|^{2}\mathrm{\leavevmode\nobreak\ d}\mathbf{x}\mathrm{\leavevmode\nobreak\ d}t,$ (14)
$w(t)=\exp\left(-\epsilon\int_{0}^{t}\int_{\Omega}\left|\mathcal{R}\left[\mathbf{u}_{\theta}\right](\tau,\mathbf{x})\right|^{2}\mathrm{\leavevmode\nobreak\ d}\mathbf{x}\mathrm{\leavevmode\nobreak\ d}\tau\right).$ (15)
@@
[…]
@@
To further enhance model performance and reduce optimization difficulty, we employ a time-marching strategy that decomposes the temporal domain into sequential windows. Rather than training a PINN model on the entire time domain simultaneously, we solve the problem progressively by training a separate PINN model for each consecutive time interval $[t_{i},t_{i+1}]$ for $i=0,1,\ldots,N-1$, where $t_{0}=0$ and $t_{N}=T$.
For each time window, we implement transfer learning by initializing the PINN parameters using the parameter values of the converged PINN model from the previous time interval [31, 32]. Specifically, the network parameters $\theta^{(i+1)}$ for time window $[t_{i+1},t_{i+2}]$ are initialized with the optimized parameters $\theta^{(i)}_{*}$ from the preceding window $[t_{i},t_{i+1}]$. The initial condition for each new time window is obtained from the PINN solution at the final time of the previous interval, ensuring continuity across temporal segments. Our experiments demonstrate that this time-marching strategy with transfer learning leads to significantly improved loss convergence and solution accuracy compared to training over the entire temporal domain simultaneously from scratch.
@@ Appendix B Metrics
Relative $\text{Relative }L^{2}\text{ Error}=\frac{\|\tilde{\mathbf{u}}-\mathbf{u}\|_{2}}{\|\mathbf{u}\|_{2}},$ (21)
@@
[…]
@@
[…]
@@
Mean velocity profile.
The mean velocity profiles for the velocity component are computed by averaging over the streamwise ($x$) and spanwise ($z$) directions. For turbulent channel flow, the profiles are defined as
[…]
where $u_{i}\in\{u,v,w\}$ represents the streamwise, wall-normal, and spanwise velocity components, respectively. These profiles provide insight into the large-scale structure of the flow and how well the model captures the mean dynamics.
Root mean square (RMS) velocity fluctuations.
The RMS velocity fluctuations for each component are computed as
$u_{i,\text{rms}}(y)=\sqrt{\langle u_{i}^{\prime 2}\rangle},\quad u_{i}^{\prime}=u_{i}-\langle u_{i}\rangle,$ (27)
where the averaging is also taken over the $x$ and $z$ directions:
[…]
These metrics quantify the turbulence intensity and directional anisotropy of the flow.
@@
$\langle u^{\prime}_{i}u^{\prime}_{j}\rangle=\langle(u_{i}-\langle u_{i}\rangle)(u_{j}-\langle u_{j}\rangle)\rangle.$ (29)
In turbulent flows such as channel flow, the $\langle u^{\prime}v^{\prime}\rangle$ component is especially important as it represents shear stress due to turbulence. Accurate modeling of Reynolds stresses is crucial for reproducing the correct momentum balance.
@@ Appendix C Training protocols
The model is trained by minimizing the composite loss over $10^{5}$ iterations using the SOAP optimizer, with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and eigenspace updates every $f=2$ steps. The learning rate increases linearly to $10^{-3}$ in the first 22,000 steps, followed by an exponential decay with a factor of 0.9 every 2,000 steps.
@@
We apply this multi-stage procedure within each time window up to $T=8$ for this problem. Specifically, we use a two-stage setup with prefactors $\epsilon_{0}=1$ and $\epsilon_{1}=0.001$. However, we find it effective only when the first-stage network already achieves sufficiently low PDE residuals (e.g., $<10^{-8}$). In cases where the initial PDE residual remains large, subsequent stages fail to effectively reduce the error. For the Kolmogorov flow and turbulent channel flow, achieving this level of accuracy in the initial stage proves difficult. As a result, the multi-stage scheme offers limited improvement for these cases and is not employed.

### v2

@@ Abstract
Turbulent fluid flows are among the most computationally demanding problems in science, requiring enormous computational resources that become prohibitive at high flow speeds. Physics-informed neural networks (PINNs) represent a radically different approach that trains neural networks directly from physical equations rather than data, offering the potential for continuous, mesh-free solutions. Here we show that appropriately designed PINNs can successfully simulate fully turbulent flows in both two and three dimensions, directly learning solutions to the fundamental fluid equations without traditional computational grids or training data. Our approach combines several algorithmic innovations including adaptive network architectures, causal training, and advanced optimization methods to overcome the inherent challenges of learning chaotic dynamics. Through rigorous validation on challenging turbulent problems, we demonstrate that PINNs accurately reproduce key flow statistics including energy spectra, kinetic energy, enstrophy, and Reynolds stresses. Our results demonstrate that neural equation solvers can handle complex chaotic systems, opening new possibilities for continuous turbulence modeling that transcends traditional computational limitations.
@@ Results
on a triply periodic domain $0,2\pi]^{3}$. The flow evolves through vortex stretching, roll-up, and eventual breakdown, leading to a fully developed turbulent state. Our objective is to simulate the flow evolution up to final time $T=10$.
@@
Building on our successful simulation of the TGV problem, we proceed to assess the framework’s performance in three-dimensional turbulent channel flow. This prototypical wall-bounded system represents a significantly more stringent test for numerical methods. Unlike the periodic or isotropic flows previously considered, channel flow is characterized by sharp gradients near solid boundaries, inherent anisotropy, and nontrivial momentum transport. For a rigorous evaluation of predicted turbulence statistics, we specifically focus on the widely adopted benchmark case at a friction Reynolds number of $\mathrm{Re}_{\tau}=395$, for which extensive DNS statistics data are available [42, 43].
@@
We simulate incompressible channel flow within a computational domain adopting the standard configuration: $L_{x}=2\pi$ in the streamwise direction, $L_{z}=\pi$ in the spanwise direction, and a wall-normal height of $L_{y}=2$. No-slip boundary conditions are applied at the channel walls ($y=\pm 1$), with periodicity enforced in the streamwise ($x$) and spanwise ($z$) directions. We set $u_{\tau}=0.09875$ and the flow is driven by a constant pressure gradient of $dp/dx=u_{\tau}^{2}$, with the kinematic viscosity set to $\nu=0.00025$. Importantly, our PINN simulation uses a snapshot of a turbulent DNS field as the initial condition. Our goal is to examine if PINNs can maintain and accurately reproduce the established turbulence statistics for this system over the time interval $t\in[0,10]$.
@@
Figure 4 shows the framework’s performance on the turbulent channel flow benchmark, compared against the reference DNS data[42, 43, 44]. Figure 4a displays a representative instantaneous snapshot of the predicted velocity norm, which clearly captures the characteristic near-wall streaks and large-scale turbulent structures inherent to wall-bounded turbulence. Beyond this qualitative assessment, we quantify the learned statistics against reference DNS data from [42, 43, 44].
Figure 4b shows the mean streamwise and spanwise velocity profiles together with the mean pressure distribution, all normalized in wall units. The PINN predictions closely follow the DNS data across the viscous sublayer, buffer layer, and logarithmic region, demonstrating the model’s ability to recover canonical mean flow behavior. Some discrepancies arises in the spanwise velocity component: because its magnitude is relatively small, the PINN struggles to capture it with the same accuracy as the streamwise counterpart. Figure 4c presents the Reynolds stress components $\overline{u^{\prime}u^{\prime}}$$^{+}$, $\overline{v^{\prime}v^{\prime}}$$^{+}$, $\overline{w^{\prime}w^{\prime}}$$^{+}$, and $\overline{u^{\prime}v^{\prime}}$$^{+}$, normalized by the friction velocity $u_{\tau}$. Despite minor deviations, the predicted profiles show good agreement with DNS data in both magnitude and wall-normal distribution, successfully capturing the anisotropy of turbulent fluctuations.
Further validation is provided in Figures 1518, which present the mean profiles of the RMS vorticity components and the energy spectra of the velocity components at different wall-normal locations $y^{+}$. The vorticity statistics confirm that the PINN accurately captures the near-wall peak and the wall-normal decay of all vorticity components, in close agreement with DNS data. The energy spectra further demonstrate that the predicted velocity fields reproduce the correct distribution of turbulent kinetic energy across scales, although the spectral amplitudes are slightly lower than the DNS reference. These results highlight that the PINN framework is capable of resolving both near-wall dynamics and multi-scale energy transfer, while still exhibiting a mild underprediction in spectral energy content.
@@
Figure 4: Turbulent channel flow ($\text{Re}_{\tau}\sim 395$). (a) Instantaneous snapshot of the predicted velocity norm, illustrating characteristic near-wall streaks and large-scale turbulent structures. (b) Comparison of the mean streamwise and spanwise velocity profile and pressrure profile between PINN predictions and DNS data, normalized in wall units. (c) Comparison of Reynolds stress components ($\overline{u^{\prime}u^{\prime}}^{+}$, $\overline{v^{\prime}v^{\prime}}^{+}$, $\overline{w^{\prime}w^{\prime}}^{+}$, and $\overline{u^{\prime}v^{\prime}}^{+}$) between PINN predictions and DNS data, normalized by the respective friction velocity $u_{\tau}$.
@@ Methods
$\mathcal{L}(\theta)=\underbrace{\int_{\Omega}\left|\mathcal{I}\left[\mathbf{u}_{\theta}\right](\mathbf{x})\right|^{2}\mathrm{\penalty 10000\ d}\mathbf{x}}_{\mathcal{L}_{\mathrm{ic}}(\theta)}+\underbrace{\int_{0}^{T}\int_{\partial\Omega}\left|\mathcal{B}\left[\mathbf{u}_{\theta}\right](t,\mathbf{x})\right|^{2}\mathrm{\penalty 10000\ d}\mathbf{x}\mathrm{\penalty 10000\ d}t}_{\mathcal{L}_{\mathrm{bc}}(\theta)}+\underbrace{\int_{0}^{T}\int_{\Omega}\left|\mathcal{R}\left[\mathbf{u}_{\theta}\right](t,\mathbf{x})\right|^{2}\mathrm{\penalty 10000\ d}\mathbf{x}\mathrm{\penalty 10000\ d}t}_{\mathcal{L}_{\mathrm{pde}}(\theta)},$ (6)
[…]
@@
$\mathcal{L}_{\text{pde }}(\theta)=\int_{0}^{T}w(t)\int_{\Omega}\left|\mathcal{R}\left[\mathbf{u}_{\theta}\right](t,\mathbf{x})\right|^{2}\mathrm{\penalty 10000\ d}\mathbf{x}\mathrm{\penalty 10000\ d}t,$ (14)
$w(t)=\exp\left(-\epsilon\int_{0}^{t}\int_{\Omega}\left|\mathcal{R}\left[\mathbf{u}_{\theta}\right](\tau,\mathbf{x})\right|^{2}\mathrm{\penalty 10000\ d}\mathbf{x}\mathrm{\penalty 10000\ d}\tau\right).$ (15)
@@
where $\square\in\{\mathrm{ic},\mathrm{bc},\mathrm{pde}\}$ and $\|\cdot\|$ denotes the $L^{2}$ norm. Then we obtain
@@
To further enhance model performance and reduce optimization difficulty, we employ a time-marching strategy that decomposes the temporal domain into sequential windows. Rather than training a PINN model on the entire time domain simultaneously, we solve the problem progressively by training a separate PINN model for each consecutive time interval $t_{i},t_{i+1}]$ for $i=0,1,\ldots,N-1$, where $t_{0}=0$ and $t_{N}=T$.
For each time window, we implement transfer learning by initializing the PINN parameters using the parameter values of the converged PINN model from the previous time interval [31, 32]. Specifically, the network parameters $\theta^{(i+1)}$ for time window $t_{i+1},t_{i+2}]$ are initialized with the optimized parameters $\theta^{(i)}_{*}$ from the preceding window $t_{i},t_{i+1}]$. The initial condition for each new time window is obtained from the PINN solution at the final time of the previous interval, ensuring continuity across temporal segments. Our experiments demonstrate that this time-marching strategy with transfer learning leads to significantly improved loss convergence and solution accuracy compared to training over the entire temporal domain simultaneously from scratch.
@@ Appendix A Nomenclature
$\omega_{x}$, $\omega_{y}$, $\omega_{z}$ Vorticity fluctuations (wall units)
$\omega_{x}^{rms}$, $\omega_{y}^{rms}$, $\omega_{z}^{rms}$ Root-mean-square (rms) of vorticity fluctuations(wall units)
@@ Appendix B Metrics
$\text{Relative }L^{2}\text{ Error}=\frac{\|\tilde{\mathbf{u}}-\mathbf{u}\|_{2}}{\|\mathbf{u}\|_{2}},$ (21)
@@
$\text{KE}=\frac{1}{2}\int_{\Omega}\|\mathbf{u}(\mathbf{x})\|^{2}\,d\mathbf{x},$ (22)
@@
$\text{Enstrophy}=\frac{1}{2}\int_{\Omega}\|\nabla\times\mathbf{u}(\mathbf{x})\|^{2}\,d\mathbf{x}.$ (24)
@@
Mean velocity and pressure profiles (space–time average).
For statistically steady turbulent channel flow, the mean velocity and pressure profiles are obtained by averaging over the streamwise ($x$), spanwise ($z$), and time ($t$) directions:
$\overline{u}_{i}(y)=\frac{1}{T\,L_{x}L_{z}}\int_{t_{0}}^{t_{0}+T}\int_{0}^{L_{x}}\int_{0}^{L_{z}}u_{i}(x,y,z,t)\,dz\,dx\,dt,\qquad u_{i}\in\{u,v,w\},$ (26)
and
$\overline{p}(y)=\frac{1}{T\,L_{x}L_{z}}\int_{t_{0}}^{t_{0}+T}\int_{0}^{L_{x}}\int_{0}^{L_{z}}p(x,y,z,t)\,dz\,dx\,dt.$ (27)
These Reynolds (space–time) averages yield profiles that depend only on the wall-normal coordinate $y$, providing a robust measure of how well the model captures the mean dynamics.
To facilitate comparison with reference DNS data, the mean quantities are expressed in wall units:
$U^{+}(y^{+})=\frac{\overline{u}(y)}{u_{\tau}},\qquad W^{+}(y^{+})=\frac{\overline{w}(y)}{u_{\tau}},\qquad P^{+}(y^{+})=\frac{\overline{p}(y)}{\rho u_{\tau}^{2}},$
where
$y^{+}=\frac{yu_{\tau}}{\nu},\qquad u_{\tau}=\sqrt{\tau_{w}/\rho},$
with $\tau_{w}$ the mean wall shear stress and $\nu$ the kinematic viscosity.
Root mean square (RMS) velocity and vorticity fluctuations.
The RMS velocity fluctuations for each component are defined as
$u_{i}^{\text{rms}}(y)=\sqrt{\overline{u_{i}^{\prime 2}}},\qquad u_{i}^{\prime}=u_{i}-\overline{u_{i}},$ (28)
where the averaging is performed over the streamwise ($x$), spanwise ($z$), and time ($t$) directions:
$\overline{u_{i}^{\prime 2}}(y)=\frac{1}{TL_{x}L_{z}}\int_{t_{0}}^{t_{0}+T}\int_{0}^{L_{x}}\int_{0}^{L_{z}}\left(u_{i}(x,y,z,t)-\overline{u_{i}}(y)\right)^{2}\,dz\,dx\,dt.$ (29)
These profiles quantify the turbulence intensity and directional anisotropy of the flow.
Similarly, the RMS vorticity fluctuations are given by
$\omega_{i}^{\text{rms}}(y)=\sqrt{\overline{\omega_{i}^{2}}},$ (30)
where $\omega_{i}\in\{\omega_{x},\omega_{y},\omega_{z}\}$ are the components of the vorticity field, and
$\overline{\omega_{i}^{2}}(y)=\frac{1}{TL_{x}L_{z}}\int_{t_{0}}^{t_{0}+T}\int_{0}^{L_{x}}\int_{0}^{L_{z}}\omega_{i}^{2}(x,y,z,t)\,dz\,dx\,dt.$ (31)
These vorticity-based statistics highlight the strength of small-scale rotational structures and provide complementary insight into the near-wall dynamics.
@@
$\overline{u^{\prime}_{i}u^{\prime}_{j}}=\overline{(u_{i}-\overline{u_{i}})(u_{j}-\overline{u_{j}})}.$ (32)
In turbulent flows such as channel flow, the $\overline{u^{\prime}v^{\prime}}$ component is especially important as it represents shear stress due to turbulence. Accurate modeling of Reynolds stresses is crucial for reproducing the correct momentum balance.
Velocity energy spectrum.
The energy spectrum for each velocity component at different wall-normal locations characterizes the distribution of kinetic energy across wavenumbers. For the velocity component $u_{i}$, the one-dimensional energy spectrum in the $x_{i}$-direction is computed as
$E_{u_{i}u_{i}}(k_{i},y)=\left\langle\sum_{k_{j}}\hat{u}_{i}(k_{i},k_{j},y),\hat{u}_{i}^{*}(k_{i},k_{j},y)\right\rangle,$ (33)
where $k_{i}$ is the streamwise ($i=1$) or spanwise ($i=3$) wavenumber, $k_{j}$ is the complementary horizontal wavenumber, $\hat{u}_{i}$ denotes the Fourier coefficient of $u_{i}$, and $(\cdot)^{*}$ indicates the complex conjugate. The angle brackets $\left\langle\cdot\right\rangle$ denote ensemble averaging over multiple samples to obtain a statistically converged spectrum.
@@ Appendix C Training protocols
The model is trained by minimizing the composite loss over $10^{5}$ iterations using the SOAP optimizer, with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and eigenspace updates every $f=2$ steps. The learning rate increases linearly to $10^{-3}$ in the first 22,000 steps, followed by an exponential decay with a factor of 0.9 every 2,000 steps. To further accelerate convergence, we employ the schedule-free method of [52] with momentum $\beta=0.9$, together with gradient clipping (global norm of 1) to ensure training stability.
@@
We apply this multi-stage procedure within each time window up to $T=8$ for this problem. Specifically, we use a two-stage setup with prefactors $\epsilon_{0}=1$ and $\epsilon_{1}=0.001$. However, we find it effective only when the first-stage network already achieves sufficiently low PDE residuals (e.g., $10^{-8}$). In cases where the initial PDE residual remains large, subsequent stages fail to effectively reduce the error. For the Kolmogorov flow and turbulent channel flow, achieving this level of accuracy in the initial stage proves difficult. As a result, the multi-stage scheme offers limited improvement for these cases and is not employed.